OpenStudy (alison123):

prove that lim as x approaches 3 of (x^2+x-4)=8 using epsilon and delta. explain steps please

OpenStudy (anonymous):

\[\begin{align*}\left|\left(x^2+x-4\right)-8\right|&=\left|x^2+x-12\right|\\ &=\left|(x-3)(x+4)\right|\\ &=|x-3||x+4| \end{align*}\] Suppose we agree to set \(\delta\le1\), then \[\begin{align*} |x-3|<1~~\implies&~~-1<x-3<1\\ &~~~~~~~6<x+4<8\\ &~~~~~~~~~~~~|x+4|<8 \end{align*}\] Then we have \[\begin{align*}\left|\left(x^2+x-4\right)-8\right|&=|x-3||x+4|\\ &<8|x-3| \end{align*}\] So for this to be less than \(\epsilon\), we can set \(\delta=\min\left\{1,\dfrac{\epsilon}{8}\right\}\)